# 1]organization=Scuola Normale Superiore and INFN, addressline=Piazza dei Cavalieri 7, city=Pisa, postcode=56126, state=PI, country=Italy

### v2

@@ Abstract
We explore the phenomenology of the “Dark Walker”—an $\text{SU}(3)$ theory with eight flavors of massless fundamental fermions in the dark sector. During inflation, its walking dynamics generate primordial non-Gaussianities through the exchange of unparticles, while accounting for the current dark matter relic abundance if the resulting massive pseudoscalars behave as strongly interacting massive particles (SIMPs). This provides a simple yet concrete example of a strongly coupled sector during inflation.
@@
In this Letter, we consider a dark sector that interacts with the Standard Model (SM) through the Higgs portal and becomes strongly coupled during inflation. After inflation, this dark sector contributes to today’s relic abundance of dark matter (DM). As a simple illustrative example, we study what we call the “Dark Walker”—a scenario that has the potential to account simultaneously for non-Gaussianities and the observed DM relic abundance.
Setup.—To investigate a strongly coupled dark sector during inflation, it is useful to classify different scenarios according to the mass gap: for example, $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$, $\text{M}_{\text{Gap}}\sim H_{\text{inf}}$ and $\text{M}_{\text{Gap}}\gg H_{\text{inf}}$. In this work, the Dark Walker model falls into the $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$ regime, where the dark sector is well approximated by a conformal field theory. Such a conformal sector may contain operators with large anomalous dimensions, which are called unparticles, as introduced in [6, 7], etc. An extended notion of unparticles with a mass gap has also been discussed in [8], but lies beyond the scope of this work. Gapless unparticles, however, cannot serve as dark matter, since their behavior resembles radiation rather than particles. Thus, the Dark Walker has to experience a phase transition, forming bound states. During inflation, the model remains near the infrared fixed point (IRFP) long enough to leave imprints in the primordial non-Gaussianity. This behavior exemplifies a form of “walking dynamics”, which gives rise to the name “Dark Walker”.
Let us consider a non-Abelian gauge theory in the dark sector (DS), which is an $\text{SU}(N_{c})_{D}$ theory with $N_{f}$ flavors of massless Dirac fermions. At the start of inflation, all modes are very blue-shifted and the effect of the de Sitter curvature in the $\beta$-function will be small, so we can approximately treat the background as flat space. The Lagrangian is
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors of massless Dirac fermions. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with MFP.
Into inflation.—In this work, we choose the value of $H_{\text{inf}}$ to be one power of ten times smaller than the upper bound value from [14], which is $H_{\text{inf}}\sim 10^{-1}\times 2.5\times 10^{-5}M_{\text{pl}}\sim 6.0\times 10^{12}\text{ GeV}$. In the context of cosmology, it is convenient to use the number of $e$-folds $N$ to measure the ratio between scales. There are $\sim 13$ $e$-folds between the Planck scale and the Hubble scale at the start of inflation.
If we start from a weak gauge coupling in the UV, we can estimate the number of $e$-folds needed to arrive near the merged fixed point (MFP) using
@@
$\Rightarrow\ N_{\text{cw}}&=\frac{8\pi^{2}}{\beta_{0}}\left(\frac{1}{g_{0}^{2}}-\frac{1}{g_{c}^{2}}\right)\\ &+\frac{\beta_{1}}{2\beta_{0}^{2}}\log\left(\frac{g_{c}^{2}}{g_{0}^{2}}\frac{16\pi^{2}\beta_{0}-g_{0}^{2}\beta_{1}}{16\pi^{2}\beta_{0}-g_{c}^{2}\beta_{1}}\right).\end{aligned}$ (6)
We will take $g_{c}^{2}$ to be the lattice result of [13], where $g_{\text{GF}}^{2}\approx 25$ will give us $g_{c}^{2}\lesssim 12$, with reference to [15]. Surely, this one-loop matching in [15] will break up for strong couplings, but all we need is an upper bound for the number of $e$-folds needed.
Since $g_{c}/4\pi\lesssim 1$, (6) is enough for an estimate. Based on the number of $e$-folds between $M_{\text{pl}}$ and $H_{\text{inf}}$, to prepare the unparticles, we expect less than 10 $e$-folds to achieve the MFP. Taking $g_{0}=2$, the number of $e$-folds needed and the resulting energy scale for forming the conformal window will be
$&N_{\text{cw}}\approx 4.0,\\ \Rightarrow\ &\Lambda_{\text{cw}}=e^{-4.0}M_{\text{pl}}\approx 4.4\times 10^{16}\text{ GeV}.\end{aligned}$ (7)
Once our model reaches the MFP, the gauge theory (1) becomes a conformal theory, and we cannot use this Lagrangian. In general, it is not easy to derive the operator spectrum and the operator product expansion (OPE) coefficients between operators from the UV theory (1). Nevertheless, we can focus on a few operators that we expect to see from the UV theory. An important one is the primary scalar operator $\mathcal{O}_{\psi}\equiv\bar{\psi}\psi$, as the anomalous dimension $\gamma_{m}$ of $\mathcal{O}_{\psi}$ will be greater than 1 at the MFP, studied in [16]. This property will trigger “walking” and drag the Dark Walker into the gapped phase, as we will see below. Another possible primary scalar operator is $\mathcal{O}_{F}\equiv\text{Tr}F^{2}$, where the scaling dimension $\Delta_{F}\lesssim 4$. The explicit value can only be found from lattice simulations. The operator $\mathcal{O}_{\psi}$ has the scaling dimension $\Delta_{\psi}=3-\gamma_{m}\approx 1.9$ from [16], and we assume $\Delta_{F}\approx 3.9$ for $\mathcal{O}_{F}$.
Using the language of the effective field theory (EFT) of inflation [17, 18], we can construct the interactions between the scalar primary operators $\mathcal{O}_{\Delta}$ and the scalar perturbation $\pi$ as given in [1, 5]:
@@
It is worth mentioning that now we should regard the unparticles $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ as composite particles, which means, for example, $\dot{\pi}\mathcal{O}_{\psi}$ is a linear mixing instead of a Yukawa coupling. The dimension of the linear mixing $\dot{\pi}\mathcal{O}_{\psi}$ is smaller than four, making this linear mixing relevant. This interaction will walk the theory from the MFP into the gapped phase, while a small relevant interaction will not affect the overall dynamics of inflation.
Since now the power of the coupling constant depends on $\Delta_{\psi}$ and $\Delta_{F}$, we need the effective coupling constant $\mu_{\psi}$ to be much smaller than $H_{\text{inf}}$ and $\mu_{F}$ to be much larger than $H_{\text{inf}}$, in order to have a sufficient perturbative description. In a slow-roll model, we can take
$\mu_{\psi}$ $=|\dot{H}_{\text{inf}}|/\Lambda_{\text{cw}}\approx 7.9\times 10^{6}\text{ GeV}\ll H_{\text{inf}},$ (9)
$\mu_{F}$ $=\dot{\phi}/\Lambda_{\text{cw}}\approx 4.6\times 10^{13}\text{ GeV}\gg H_{\text{inf}},$ (10)
which is good enough for perturbative expansion. In addition, assuming that $\langle\mathcal{O}_{\Delta}^{3}\rangle$ is negligible, the bispectrum is determined by the quadratic derivative coupling $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}$, as discussed in [5].
@@
$f_{\text{NL}}\equiv\frac{5}{18}\frac{\mathcal{B}_{\zeta}(k,k,k)}{\Delta_{\zeta}^{4}(k)},$ (11)
where $\mathcal{B}_{\zeta}$ is the bispectrum of curvature perturbation $\zeta$, and $\Delta_{\zeta}^{2}=2.1\times 10^{-9}$ is the dimensionless power spectrum of $\zeta$ from [19]. The $f_{\text{NL}}$ using the language of the EFT of inflation can be written as
$f_{\text{NL}}=\frac{5}{18}\frac{1}{\Delta_{\zeta}^{4}}\left(\frac{1}{2}\mu^{2-\Delta}M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}\right)^{2}H_{\text{inf}}^{2\Delta-1}$
×$\times\frac{1}{M_{\text{pl}}^{2}|{\dot{H}_{\text{inf}}}|}\left(-\frac{H_{\text{inf}}}{\dot{\phi}}\right)^{3}\times\frac{(2\pi)^{4}}{16k^{6}}S(k,k,k).$ (12)
@@
Taking the slow-roll parameter $\epsilon$ to be $0.0097$ according to [14], we find that $f_{\text{NL},\psi}\approx 16$ and $f_{\text{NL},F}\approx 3.5\times 10^{-3}$, which are desirable because they are around the center of the latest observation $f_{\text{NL}}=6\pm 46$, as shown in [20]. Notice that the choice of $H_{\text{inf}}$ will significantly affect the value of $f_{\text{NL}}$. Figure 2 shows the shape functions with $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ exchanges. Both of them are near-orthogonal, while the squeezed limit behaviors degenerate with the equilateral non-Gaussianity. Only the full shapes of the bispectra can break this degeneracy, as emphasized in [5].
@@
Figure 2: Shape functions for $\Delta=1.9$ and $\Delta=3.9$. We rescale the magnitudes by $10^{2}$ for $\Delta=1.9$ and $10^{3}$ for $\Delta=3.9$ to set the plot value within order 1.

### v3

@@ Abstract
We explore the phenomenology of the “Dark Walker”—an $\text{SU}(3)$ theory with eight flavors of massless fundamental fermions in the dark sector. During inflation, its walking dynamics generate primordial non-Gaussianities through the exchange of unparticles, while accounting for the current dark matter relic abundance if we consider freeze-in of Dark Walker coupled to the Standard Model through the Higgs portal. This provides a simple yet predictive example linking strongly coupled inflationary dynamics to present-day dark matter.
@@
In this Letter, we consider a dark sector (DS) that interacts with the Standard Model (SM) through the Higgs portal and becomes strongly coupled during inflation, leaving imprints in the primordial fluctuations. After inflation, this DS contributes to today’s relic abundance of dark matter (DM). We aim to explore the interplay between primordial non-Gaussianity and DM model building. As a simple illustrative example, we study what we call the “Dark Walker”—a scenario that has the potential to account simultaneously for primordial non-Gaussianity $f_{\text{NL}}$ value at $O(1)$ and the observed DM relic abundance.
Setup.—To investigate a strongly coupled DS during inflation, it is useful to classify different scenarios according to the mass gap: for example, $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$, $\text{M}_{\text{Gap}}\sim H_{\text{inf}}$ and $\text{M}_{\text{Gap}}\gg H_{\text{inf}}$. In this work, the Dark Walker model falls into the $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$ regime, where the DS is well approximated by a conformal field theory. Such a conformal sector may contain operators with large anomalous dimensions, which are called unparticles, as introduced in [6, 7], etc. An extended notion of unparticles with a mass gap has also been discussed in [8], but lies beyond the scope of this work. Gapless unparticles, however, cannot serve as DM, since their behavior resembles radiation rather than particles. Thus, the Dark Walker has to experience a phase transition, forming gapped states. During inflation, the model remains near the infrared fixed point (IRFP) long enough to leave imprints in the primordial non-Gaussianity. This behavior exemplifies a form of “walking dynamics”, which gives rise to the name “Dark Walker”.
Let us consider a non-Abelian gauge theory in the DS, which is an $\text{SU}(N_{c})_{D}$ theory with $N_{f}$ flavors of massless Dirac fermions. At the start of inflation, all modes are very blue-shifted and the effect of the de Sitter curvature in the $\beta$-function will be small, so we can approximately treat the background as flat space. The Lagrangian is
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors of massless Dirac fermions. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with a merged fixed point (MFP). This kind of mass generation is called the symmetric mass generation (SMG), a phenomenon where the fermions acquire mass gaps while the chiral symmetry remains unbroken. For a recent review, the readers can refer to [14].
During Inflation.—In this work, we choose the value of $H_{\text{inf}}$ to be the current upper bound value from [15], which is $H_{\text{inf}}\sim 2.5\times 10^{-5}M_{\text{pl}}\sim 6.0\times 10^{13}\text{ GeV}$. In the context of cosmology, it is convenient to use the number of $e$-folds NN to measure the ratio between scales. There are $\sim 10$ $e$-folds between the Planck mass scale and the Hubble scale during inflation. Once we assume the inflation starts from the $M_{\text{pl}}$, the working energy scale during inflation flows from $M_{\text{pl}}$ to $H_{\text{inf}}$, and stays at $H_{\text{inf}}$ until the end of inflation.
If we start from a weak gauge coupling in the UV, we can estimate the number of $e$-folds needed to arrive near the MFP using
@@
$\begin{aligned} \Rightarrow\ N_{\text{cw}}&=\frac{8\pi^{2}}{\beta_{0}}\left(\frac{1}{g_{0}^{2}}-\frac{1}{g_{c}^{2}}\right)\\ &+\frac{\beta_{1}}{2\beta_{0}^{2}}\log\left(\frac{g_{c}^{2}}{g_{0}^{2}}\frac{16\pi^{2}\beta_{0}-g_{0}^{2}\beta_{1}}{16\pi^{2}\beta_{0}-g_{c}^{2}\beta_{1}}\right).\end{aligned}$ (6)
We will take the critical gauge coupling $g_{c}^{2}$ to be the lattice result of [13]. The critical gradient-flow coupling $g_{\text{GF}}^{2}\approx 25$ will give us $g_{c}^{2}\lesssim 12$, with reference to [16]. Surely, this one-loop matching in [16] will break up for strong couplings, but all we need is an upper bound for the number of $e$-folds needed. We will restrict the ranges of $N_{\text{cw}}$ and the initial value of the gauge coupling $g_{0}$ after constructing the effective interactions.
Once our model reaches the MFP, the gauge theory (1) becomes a conformal theory, and we cannot use this Lagrangian. Generally, it is not easy to derive the operator spectrum and the operator product expansion (OPE) coefficients between operators from the UV theory (1). Nevertheless, we can focus on a few operators that we expect to see from the UV theory. An important one is the fermion bilinear operator $\mathcal{O}_{\psi}\equiv\bar{\psi}\psi$, as the anomalous dimension $\gamma_{m}$ of $\mathcal{O}_{\psi}$ will be greater than 1 near the MFP, studied in [17]. This property will trigger “walking” and drag the Dark Walker into the gapped phase, as we will see below. Another possible primary scalar operator is $\mathcal{O}_{F}\equiv\text{Tr}F^{2}$, where the scaling dimension $\Delta_{F}\lesssim 4$. The explicit value can only be found from lattice simulations. The operator $\mathcal{O}_{\psi}$ has the scaling dimension $\Delta_{\psi}=3-\gamma_{m}\approx 1.8$ from [17], and we assume $\Delta_{F}\approx 3.9$ for $\mathcal{O}_{F}$.
Using the language of effective field theory (EFT) of inflation [18, 19], we can construct the linear interactions between the scalar primary operators $\mathcal{O}_{\Delta}$ and the scalar perturbation $\pi$ as given in [1, 5]:
@@
It is worth mentioning that now we should regard the unparticles $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ as composite particles, which means, for example, $\dot{\pi}\mathcal{O}_{\psi}$ is a linear mixing instead of a Yukawa coupling. The dimension of the linear mixing $\dot{\pi}\mathcal{O}_{\psi}$ is less than four, making this linear mixing relevant. This interaction will walk the theory from the MFP into the gapped phase, while a small relevant interaction will not affect the overall dynamics of inflation.
Since now the power of the coupling constant depends on $\Delta_{\psi}$ and $\Delta_{F}$, we need the effective coupling constant $\mu_{\psi}$ to be smaller than $H_{\text{inf}}$ and $\mu_{F}$ to be larger than $H_{\text{inf}}$, in order to have a sufficient perturbative description. In a slow-roll inflation model, it is natural to take $\mu^{2-\Delta}=\dot{\phi}/\Lambda_{\text{cw}}^{\Delta}$, where $\Lambda_{\text{cw}}$ is the energy scale to form the unparticles. Since $g_{c}/4\pi\lesssim 1$, (6) is enough for an estimate. In order to have a sufficient separation of scales to perform perturbative analysis, we require the lower bound of $\Lambda_{\text{cw}}$ to be $\Lambda_{\text{cw}}^{\Delta}\gtrsim\dot{\phi}H_{\text{inf}}^{\Delta-2}$ for all $\Delta$’s we consider. Regarding all the requirements above, the number of $e$-folds needed to arrive near the MFP will be restricted to $0\lesssim N_{\text{cw}}\lesssim 5.8$, and the initial value $g_{0}$ will be limited to $1.7\lesssim g_{0}\lesssim 3.4$. Explicitly,
$7.3\times 10^{15}\text{ GeV}\lesssim\Lambda_{\text{cw}}\lesssim M_{\text{pl}}.$ (8)
In addition, assuming that $\langle\mathcal{O}_{\Delta}^{3}\rangle$ is negligible, the bispectrum is determined by the quadratic derivative coupling $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}$, as discussed in [5]. Having established the EFT of inflation with unparticles, we now turn to its cosmological signatures.
@@
$f_{\text{NL}}\equiv\frac{5}{18}\frac{(k_{1}k_{2}k_{3})^{2}}{(2\pi^{2})^{2}\Delta_{\zeta}^{4}(k)}\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle^{\prime}\big|_{k_{1}=k_{2}=k_{3}=k},$ (9)
where
$\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle\equiv(2\pi)^{3}\delta^{(3)}(\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3})\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle^{\prime}$
is the three-point correlation function of curvature perturbation $\zeta$. $\Delta_{\zeta}^{2}=2.1\times 10^{-9}$ is the dimensionless power spectrum of $\zeta$ from [20]. At linear order, the curvature perturbation $\zeta$ relates to the scalar perturbation $\pi$ via
$\zeta=-H_{\text{inf}}\ \pi.$ (10)
@@
Taking the slow-roll parameter $\epsilon$ to be $0.0097$ according to [15], we find that $1.2\times 10^{-8}\lesssim f_{\text{NL},\psi}\lesssim-13$ and $|f_{\text{NL},F}|<10^{-10}$. The full expression of $f_{\text{NL}}$ is presented in Appendix A. The $f_{\text{NL}}$ with $\mathcal{O}_{\psi}$ is desirable because it has the potential to be $O(1)$, which is around the center of the latest observation $f_{\text{NL}}=6\pm 46$ [21], while large enough to be detectable with the experiments in the near future. The $f_{\text{NL},F}$ is too small to be detected. Even the futuristic 21-cm tomography can only give a limiting strength of non-Gaussianity for observability about $f_{\text{NL}}\gtrsim 10^{-4}$ [22]. Notice that the choice of $H_{\text{inf}}$ will significantly affect the value of $f_{\text{NL}}$, as indicated in Figure 2.
Figure 2: The contour plot of $f_{\text{NL},\psi}$ at $O(1)$ for a range of $H_{\text{inf}}$ and $\Lambda_{\text{cw}}$. The magnitude of $f_{\text{NL}}$ is sensitive to the value of $H_{\text{inf}}$.
Figure 3 shows the shape functions with $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ exchanges. Both of them are near-equilateral, while the squeezed limit behaviors degenerate with the equilateral non-Gaussianity. Only the full shapes of the bispectra can break this degeneracy, as emphasized in [5].
@@
Figure 3: Shape functions for $\Delta=1.8$ and $\Delta=3.9$. We rescale the magnitudes by $10^{2}$ for $\Delta=1.8$ and $10^{3}$ for $\Delta=3.9$ to set the plot value within $O(1)$.